OpenStudy (anonymous):

A light rope is fixed at one end of a wooden clamp on the ground passes over a branch and hangs on the other side. It makes an angle 30 degrees with the ground. A man weighing 60 kg wants to climb up the rope. The wooden clamp can come out of the ground if an upward force greater than 360N is applied. Find the max ace in the upward direction with which the man can climb safely. Neglect friction at the tree branch. Take g=10m/s^2

OpenStudy (anonymous):



OpenStudy (anonymous):

sol..Let T be the tension. Upward force is Tsin30=t/2=360N this is what i dont understand dont u take normal force into account???

OpenStudy (anonymous):

the force would be the upward force-downward force...?

OpenStudy (anonymous):

I wrote the above from the solution/ T+N-mg=360 i think this is what it should be

OpenStudy (anonymous):

T sin30 =360N and then T-360= Force and then Force=ma am i missing something?

OpenStudy (anonymous):

sorry that's 600-T=force

OpenStudy (anonymous):

this is what iam asking why is Tsin30=360 and not T+Normal force-mg=360 ?

OpenStudy (anonymous):

Normal force?

OpenStudy (anonymous):

im talking about the clamp

OpenStudy (anonymous):

oh, well actually as you can see in your diagram, the upward tension in the rope comes out to be the same if you take it from the clamp or if you take it from where the person is hanging down...the TENSION in the rope means the same thing from either side of it you take....

OpenStudy (anonymous):

why is Tsin30=360(the total upward force) and not T+Normal force-mg=360 ? with clamp as system

OpenStudy (anonymous):

The tension and normal force are the same thing.

OpenStudy (anonymous):